OpenStudy (anonymous):

2.Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (mertsj):

He could go back to his original equation and replace each x with 1 and see if it makes a true sentence

OpenStudy (texaschic101):

(3x + 6)/8 = (7x - 1)/6 (common denominator is 24) 3(3x + 6) = 4(7x - 1) -- distribute through the parenthesis 9x + 18 = 28x - 4 ===> FIRST MISTAKE 9x - 9x + 18 = 28x - 9x - 4 18 = 19x - 4 18 + 4 = 19x - 4 + 4 22 = 19x 22/19 = x

OpenStudy (texaschic101):

The answer can be checked by subbing in what you got for x, back into the equation to see if both sides equal.

OpenStudy (anonymous):

I got 1.15